# Rainbow PO: A Unified Framework for Combining Improvements in Preference Optimization

### v1

@@ Abstract
Recently, numerous preference optimization algorithms have been introduced as extensions to the Direct Preference Optimization (DPO) family. While these methods have successfully aligned models with human preferences, there is a lack of understanding regarding the contributions of their additional components. Moreover, fair and consistent comparisons are scarce, making it difficult to discern which components genuinely enhance downstream performance. In this work, we propose RainbowPO, a unified framework that demystifies the effectiveness of existing DPO methods by categorizing their key components into seven broad directions. We integrate these components into a single cohesive objective, enhancing the performance of each individual element. Through extensive experiments, we demonstrate that RainbowPO outperforms existing DPO variants. Additionally, we provide insights to guide researchers in developing new DPO methods and assist practitioners in their implementations.
@@ 1 Introduction
Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022; Stiennon et al., 2020; Ziegler et al., 2019) has significantly contributed to the success of recently released Large Language Models (LLMs) such as InstructGPT (Ouyang et al., 2022), ChatGPT, and GPT4 (Achiam et al., 2023). However, RLHF is a complex and resource intensive process and requires training a reward model. An alternative to RLHF is Direct Preference Optimization (DPO) (Rafailov et al., 2023) that directly optimizes policies from pairwise preferences by minimizing a supervised learning loss objective, which is viewed as the maximum likelihood estimate for the reward model in RLHF. This approach allows DPO and similar other DPO variants to bypass the use of RL, resulting in faster speed of end-to-end training and better resource efficiency, while achieving comparable or superior performance to RLHF in downstream tasks such as summarization (Rafailov et al., 2023).
DPO and its success during training foundation models like LLama series (Dubey et al., 2024; Touvron et al., 2023), Mistral (Jiang et al., 2023a), has garnered significant research attention in the LLM alignment space (Winata et al., 2024; Wang et al., 2024b), leading to the development of various extensions. These include variants beyond pairwise ranking, such as KTO (Ethayarajh et al., 2023; Song et al., 2024), unified perspectives on loss parameterization, such as IPO (Azar et al., 2024) and GPO (Tang et al., 2024), distribution correction methods like RSO (Liu et al., 2023) and WPO (Zhou et al., 2024), and reference model-free alternatives, such as CPO (Xu et al., 2024), ORPO (Hong et al., 2024), and SimPO (Meng et al., 2024). Each of these DPO variants claims to outperform the original DPO in downstream task evaluations by introducing specific components, or mathematically modifying the loss objective. In the rest of the paper, we will refer to all DPO variants collectively as xPOs for simplicity.
Comparing these xPOs proposed in the literature is not always straightforward due to differences in the base model size and architecture, the alignment datasets, the experimental setup as well as the evaluation metrics. Subsequently, it becomes difficult to assess the effectiveness and choose among different xPO methods given a problem. A brute force comparison across all existing methods is prohibitively expensive and inefficient. Therefore, it is crucial that we study the performance characteristics of each proposed method in the literature by evaluating the xPOs’ performances under at least one convincing and representative setup.
Further, despite the success of the xPO family, a fundamental question remains unexplored:
@@
Surprisingly, there is still a lack of comprehensive work studying the progress in the literature and summarizing the core practical components of these methods that lead to performance improvement of the DPO objective. To demystify the reasons for their effectiveness, we hypothesize that the main benefits of these methods stem from the combination of several mathematically orthogonal effective components. In this paper, we validate our hypothesis by decomposing the xPOs and identifying these orthogonal components upon DPO. We further assess their effectiveness through downstream task evaluations, ruling out the components that do not contribute to performance improvements. Given these orthogonal identified beneficial components for preference optimization, a natural question arises:
@@
    We conduct a comprehensive study on more than 10 offline representative variants of DPOs (xPOs) from a practical aspect by analyzing their loss functions for optimization. We conclude analyze several mathematically orthogonal directions along which these methods propose to optimize over the original DPO loss, analyze the usefulness of each method theoretically and empirically, and provide comparisons under the same representative setup.
@@
    We identify and summarize seven broad components across all DPO extensions: length normalization, link function, margin / home advantage, reference policy, contextual scaling, rejection sampling optimization (RSO), and supervised fine-tuning (SFT) loss, and justify that four of them are effective through extensive hyper-parameters search, model training and evaluations. Additionally, we also propose a better way of formulating the reference policy by a linear mixing of the SFT policy and the margin, and demonstrate the advantage of this approach over using just reference policy (in DPO) or just margin (in SimPO).
@@
[…]
@@ 2 Preliminaries and Motivation
RLHF starts with fine-tuning a pre-trained large language model by supervised learning on high-quality data for some downstream tasks of interest (e.g., dialogue, summarization, etc.), to acquire a model $\pi^{\mathrm{SFT}}$. This step is referred to as the SFT phase. For instance, for training InstructGPT (Ouyang et al., 2022), GPT-3 (Brown et al., 2020) is first fine-tuned on the given input prompt distribution. The second stage of RLHF is known as reward modeling, i.e., researchers collect preferences $\mathcal{D}=(x,y^{w},y^{l})$ on the generations of fine-tuned model $\pi^{\mathrm{SFT}}$, and learns a reward model $r^{*}(x,y)$ that could represent the quality or the rating of generation $y$ with respect to prompt $x$. The final step is policy optimization on $\pi_{\mathrm{SFT}}=\pi_{\mathrm{ref}}$, by maximizing a regularized reward to obtain the optimal policy model $\pi^{*}$ through reinforcement learning:
$\max_{\theta}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi_{\theta}(y\mid x)}\left[r^{*}(x,y)\right]-\beta\mathrm{KL}\left(\pi_{\theta}(\cdot\mid x)\|\pi_{\mathrm{ref}}(\cdot\mid x)\right)\right]$ (1)
For ease of reference, we add a table of notations in Table 8 in Appendix C and more detailed description of RLHF in Appendix A.1.
@@ 2.1 Direct Preference Optimization (DPO)
One disadvantage of RLHF is that the RL step often requires substantial computational effort (e.g., to carry out PPO). The idea of DPO is to combine the reward model and RL in RLHF into a single objective, bypassing the computation in the RL step. Given the same preference pairs $\mathcal{D}=(x,y^{w},y^{l})$ utilized for reward modeling in RLHF, the DPO objective yields:
@@
where $\sigma(\cdot)$ is the sigmoid function and $\beta$ is a regularization parameter for tuning. DPO thus yields a supervised learning problem, and requires much less computation than the RL based RLHF. The objective in Equation 2 can be understood as maximizing the likelihood difference between the preference pair, making the model more likely to generate the preferred answers than unpreferred.
@@ 2.2 Motivation: Revisiting xPOs
Since DPO is proposed, there is huge interest in developing and improving DPO, leading to numerous xPOs. Different xPOs can be motivated by theoretical concerns like relaxing or extending preference distribution assumptions in IPO and Mallows DPO, human aware loss function in KTO, or from practical aspects like reference model-free alternatives, like CPO, ORPO and SimPO. We provide an non-exhaustive list in Table 7 in Appendix C for the ease of comparison.
@@
Despite different motivations, xPOs all have a main loss objective that they optimize. We thus take the loss objectives as the first class citizen, and mathematically understand the parts that are commonly adopted or differ in xPOs. Before going into detailed categorization, we want to first argue that, in existing preference optimization literature, there lacks a work in revisiting and examining the DPO variants in their objectives mathematically and comprehensively. As a consequence, some papers may have implicitly proposed some designs for improvement and even didn’t highlight it. As a motivating example, we revisit ORPO objective proposes to maximize an odd ratio difference (for an event A with probability $p$, the odds ratio is defined as $p/(1-p)$):
@@
$\mathcal{L}_{\mathrm{ORPO}}\left(\pi_{\theta}\right)=-\mathbb{E}\left[\underbrace{\log p_{\theta}(y_{w}|x)}_{\mathcal{L}_{\text{ORPO-SFT}}}\underbrace{+\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right)}_{\lambda\cdot\mathcal{L}_{\text{ORPO-PO}}}\right],$ (3)
@@
in which the expectation is for $\left(x,y_{w},y_{l}\right)\sim\mathcal{D}$, and $p_{\theta}(y|x)=\exp\left(\frac{1}{|y|}\log\pi_{\theta}(y|x)\right)$. Rewriting terms in $\mathcal{L}_{\text{PO}}$, we could derive a upper bound as (see proof in Appendix B.1):
@@
$\mathcal{L}_{\mathrm{ORPO-PO}}\leq-\mathbb{E}\log\sigma\left(\frac{1}{1-p_{\theta}(y_{l}|x)}\underbrace{\left(\frac{1}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{1}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)\right)}_{\Delta_{\theta}}\right):=\bar{\mathcal{L}}_{\text{PO}},$ (4)
@@
if assuming $\Delta_{\theta}>0$ for all $x$. The upper bound ¯$\bar{\mathcal{L}}_{\text{PO}}$ is sharp, in the sense that ¯$\bar{\mathcal{L}}_{\text{PO}}-\mathcal{L}_{\text{PO}}=\mathcal{O}(\Delta_{\theta})^{2}$; thus minimizing ORPO loss could be understood as CPO with length normalization and a contextual dependent $\beta$. Length Normalization is one of the key ideas adopted in SimPO–though proposed after ORPO. This thus calls for a comprehensive analysis of the core contributed elements in different xPOs so far, as many methods may already overlap in contributed directions without awareness, and bringing this out right away could possibly prevent repetitive work or efforts in the future.
@@
Following similar analysis of different representative XPO methods for pairwise preferences, including DPO (Rafailov et al., 2023), IPO (Azar et al., 2024), CPO (Xu et al., 2024), GPO (Tang et al., 2024), RSO (Liu et al., 2023), ODPO (Amini et al., 2024), ORPO (Hong et al., 2024), Mallows-DPO (Chen et al., 2024a), SimPO (Meng et al., 2024), we come up with seven broad categories, which is able to explain most popular DPO variants in the literature, as in Table 1, This provides a straightforward illustration of the main ideas and connections of existing methods. The meanings and details of the categories are elaborated in Section 3.
@@
Mallows-DPO $\times$ logistic $\times$ SFT $\checkmark$ $\times$ $\times$
@@ 3.1 Component Descriptions
We first explain in detail about the components we categorized, after which we propose a generic framework, RainbowPO, to combine these components.
@@
The literature has noticed a verbosity issue of DPO aligned models. To address this, one promising direction noticed in the literature is to incorporate explicit verbosity penalties, like in R-DPO (Park et al., 2024) and SimPO (Meng et al., 2024):
@@
in which $r_{\theta}(x,y)=\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$ is the implicit reward model (Rafailov et al., 2023). From an optimization perspective, maximization with respect to $r^{\text{LN}}_{\theta}(x,y)$ is equivalent to $r^{\text{LR}}_{\theta}(x,y)$ with a specific $\alpha$ (might be prompt $x$ dependent). Why length normalization could help prevent the verbosity issues can be explained through examining the gradient of the loss respectively: $\nabla_{\theta}\mathcal{L}_{\mathrm{LN-DPO}}\left(\pi_{\theta};\pi_{\text{ref}}\right)=$
$-\beta\mathbb{E}\left[\sigma\left(r^{\text{LN}}_{\theta}\left(x,y_{l}\right)-r^{\text{LN}}_{\theta}\left(x,y_{w}\right)\right)\left(\frac{1}{|y^{w}|}\nabla_{\theta}\log\pi_{\theta}\left(y_{w}\mid x\right)-\frac{1}{|y^{l}|}\nabla_{\theta}\log\pi_{\theta}\left(y_{l}\mid x\right)\right)\right],$ (6)
thus the gradient of length normalized DPO can be understood as taking a discount factor $\frac{1}{|y^{w}|}$ of the length for longer sequence. We also empirically justify the effectiveness of length normalization by comparing to the vanilla DPO trained models, and witness the consistent smaller average length, independent of the regularization constant $\beta$. See results of average length in Section 4.
@@
SLiC-HF (Zhao et al., 2023) and GPO (Tang et al., 2024) both realized that the DPO objective could be understood as taking $f$ as $\log\sigma(\cdot)$ in:
$\mathcal{L}_{\text{GPO}}=\mathbb{E}\left[f\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right],$ (7)
thus unifying DPO, IPO, SLiC (without SFT loss). We did an exclusive parameter search for DPO and IPO separately, however, according to our experiments, DPO is empirically still better than IPO in validation metric as AlpacaEval for both evaluators (Llama3-70B and GPT4-1106 preview), despite the weaker assumption of preferences in IPO.
@@
which takes the losing prompt in a home advantage when $\gamma>0$. SimPO shows the effectiveness of this margin under the reference-free setup; however, when we adopt the margin for vanilla DPO (i.e. with the reference policy) with the optimal $\beta$, we do not witness an increase of the performance when adjusting the margin, either further adopting DPO with length normalization or not. In Figure 1(a), the performance steadily decreases when increasing the margin in DPO. This questions the true explanation about the effectiveness of the margin term in SimPO. We provide the answer as understanding margin as a reference policy.
@@
DPO takes the SFT policy as the reference policy motivated by the standard RLHF pipeline. However, recently proposed methods like CPO, ORPO, and SimPO (Xu et al., 2024; Hong et al., 2024; Meng et al., 2024) all suggested a reference-free objective could yield the same or even better performance. CPO and ORPO further utilized an extra SFT loss to force regularization, while for SimPO, such regularization is not enforced. Given our prior examination that home advantage can hardly improve over DPO, we argue that the margin term in SimPO should be understood as a term for “reference policy” instead of the “home advantage”.
@@
Concretely, we could hypothesize that there exists a “good policy” $\pi_{\gamma}$ such that, for each prompt and preference pairs in the dataset, the normalized log likelihood ratio of preferred response to non-preferred response is a positive constant, which we denote as $\pi_{\gamma}$. $\pi_{\gamma}$’s normalized implicit reward model is perfect at in-distribution pairwise classification and yields $\frac{\pi_{\gamma}\left(y_{w}\mid x\right)^{1/|y^{w}|}}{\pi_{\gamma}\left(y_{l}\mid x\right)^{1/|y^{l}|}}=\exp(\gamma)$ for any $x$. If so, the loss of SimPO could be rewritten as the DPO with length normalization and a different reference policy:
$\mathcal{L}_{\text{SimPO}}$ $=-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}{\mathbb{E}}\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\beta\cdot\gamma\right),$ (9)

### v2

@@ Abstract
Recently, numerous preference optimization algorithms have been introduced as extensions to the Direct Preference Optimization (DPO) family. While these methods have successfully aligned models with human preferences, there is a lack of understanding regarding the contributions of their additional components. Moreover, fair and consistent comparisons are scarce, making it difficult to discern which components genuinely enhance downstream performance. In this work, we propose RainbowPO, a unified framework that demystifies the effectiveness of existing DPO methods by categorizing their key components into seven broad directions. We integrate these components into a single cohesive objective, enhancing the performance of each individual element. Through extensive experiments, we demonstrate that RainbowPO outperforms existing DPO variants. Additionally, we provide insights to guide researchers in developing new DPO methods and assist practitioners in their implementations. Our code is available at https://github.com/CapitalOne-Research/RainbowPO.
@@ 1 Introduction
Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022; Stiennon et al., 2020; Ziegler et al., 2019) has significantly contributed to the success of recently released Large Language Models (LLMs) such as InstructGPT (Ouyang et al., 2022), ChatGPT, and GPT4 (Achiam et al., 2023). However, RLHF is a complex and resource intensive process and requires training a reward model. An alternative to RLHF is Direct Preference Optimization (DPO) (Rafailov et al., 2023) that directly optimizes policies from pairwise preferences by minimizing a supervised learning loss objective, which is viewed as the maximum likelihood estimate for the reward model in RLHF. This approach allows DPO and other DPO variants to bypass the use of RL, resulting in faster speed of end-to-end training and better resource efficiency, while achieving comparable or superior performance to RLHF in downstream tasks such as summarization (Rafailov et al., 2023).
DPO and its success during training foundation models like Llama series (Dubey et al., 2024; Touvron et al., 2023), Mistral (Jiang et al., 2023a), has garnered significant research attention in the LLM alignment space (Winata et al., 2024; Wang et al., 2024b), leading to the development of various extensions. These include variants beyond pairwise ranking, such as Kahneman & Tversky Optimization (KTO, Ethayarajh et al. (2023)) and MallowsPO (Chen et al., 2024a), unified perspectives on loss parameterization, such as Identity Preference Optimization (IPO, Azar et al. (2024)) and Generalized Preference Optimization (GPO, Tang et al. (2024)), distribution correction methods like Rejection Sampling Optimization (RSO, Liu et al. (2023)), and reference model-free alternatives, such as Contrastive Preference Optimization (CPO, Xu et al. (2024)), Odds Ratio Preference Optimization (ORPO, Hong et al. (2024)), and Simple Preference Optimization (SimPO, Meng et al. (2024)). Each of these DPO variants claims to outperform the original DPO in downstream task evaluations by introducing specific additional components, or mathematically modifying the loss objective. In the rest of the paper, we will refer DPO variants collectively as xPOs for simplicity.
Comparing these xPOs proposed in the literature is not always straightforward due to differences in the base model size and architecture, the alignment datasets, the experimental setup as well as the evaluation metrics. Subsequently, it becomes difficult to assess the effectiveness and choose among different xPO methods given a problem. A brute force comparison across all existing methods is prohibitively expensive and inefficient. Therefore, it is crucial that we study the performance characteristics of each proposed method in the literature by evaluating the xPOs’ performances under at least one convincing and representative setup. Further, despite the success of the xPO family, a fundamental question remains unexplored:
@@
Surprisingly, there is still a lack of comprehensive work studying the progress in the literature and summarizing the core practical components of these methods that lead to improvement over DPO. To demystify the reasons for their effectiveness, we hypothesize that the main benefits of these methods stem from the combination of several mathematically orthogonal effective components. In this paper, we validate our hypothesis by decomposing the xPOs and identifying these orthogonal components upon DPO. We further assess their effectiveness through downstream task evaluations, ruling out the components that do not contribute to performance improvements. Given these orthogonal identified beneficial components for preference optimization, a natural question arises:
@@
    We conduct a comprehensive study on more than 10 offline representative variants of DPOs (xPOs) from a practical aspect by analyzing their loss functions for optimization. We conclude several mathematically orthogonal directions along which these methods propose to optimize over the original DPO loss, analyze the usefulness of each method theoretically and empirically, and provide comparisons under the same representative setup.
@@
    We identify and summarize 7 broad components across all DPO extensions: length normalization, link function, margin / home advantage, reference policy, contextual scaling, rejection sampling optimization (RSO), and supervised fine-tuning (SFT) loss, and justify that four of them are effective through extensive hyper-parameters search, model training and evaluations. Additionally, we also propose a better way of formulating the reference policy by mixing the SFT policy with a margin (see details in the reference policy of Section 3.1), and demonstrate the advantage of this approach over using just SFT policy (in DPO) or just margin (in SimPO).
@@
    Finally, we propose RainbowPO, a DPO variant that combines three essential and orthogonal components from existing xPOs. Combining other adjustment on optimization hyper-parameters, we show that our algorithms perform the best among all open-sourced algorithms when tuning Llama3-8B-Instruct. In the widely adopted LLM benchmark Alpaca-Eval2  [^1] https://github.com/tatsu-lab/alpaca_eval., RainbowPO improves Llama3-8B-Instruct from 22.92% to 51.66% for Length Controlled Win Rate (LC WR), with just access to a reward model to form the offline preference dataset and no further online sampling. We also perform an ablation study and show that all adopted elements in RainbowPO are indeed necessary to achieve the best result.
@@ 2 Preliminaries and Motivation
RLHF starts with fine-tuning a pre-trained large language model by supervised learning on high-quality data for some downstream tasks of interest (e.g., dialogue, summarization, etc.), to acquire a model $\pi^{\mathrm{SFT}}$. This step is referred to as the SFT phase. For instance, for training InstructGPT (Ouyang et al., 2022), GPT-3 (Brown et al., 2020) is first fine-tuned on the given input prompt distribution. The second stage of RLHF is known as reward modeling, i.e., researchers collect preferences $\mathcal{D}=(x,y_{w},y_{l})$ on the generations of fine-tuned model $\pi^{\mathrm{SFT}}$, and learns a reward model $r^{*}(x,y)$ that could represent the quality or the rating of generation $y$ with respect to prompt $x$. The final step is policy optimization on $\pi_{\mathrm{SFT}}=\pi_{\mathrm{ref}}$, by maximizing a regularized reward to obtain the optimal policy model $\pi^{*}$ through reinforcement learning:
$\max_{\theta}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi_{\theta}(y\mid x)}\left[r^{*}(x,y)\right]-\beta\mathrm{KL}\left(\pi_{\theta}(\cdot\mid x)\|\pi_{\mathrm{ref}}(\cdot\mid x)\right)\right],$ (1)
in which $\beta>0$ denotes the regularization constant. For ease of reference, we prvide more detailed description of RLHF in Appendix A.1, and we also add a table of notations in Table 7 in Appendix B.
@@ 2.1 Direct Preference Optimization (DPO)
One disadvantage of RLHF is that the RL step often requires substantial computational effort (e.g., to carry out PPO). The idea of DPO is to combine the reward model and RL in RLHF into a single objective, bypassing the computation in the RL step. Given the same preference pairs $\mathcal{D}=(x,y_{w},y_{l})$ utilized for reward modeling in RLHF, the DPO objective yields:
@@
where $\sigma(\cdot)$ is the sigmoid function and $\beta$ is a regularization parameter similar to the one in RLHF. DPO thus yields a supervised learning problem, and requires much less computation than the RL based RLHF. The objective in Equation 2 can be understood as maximizing the likelihood difference between the preference pairs under the policy, encouraging the model to more likely generate the preferred answers than non-preferred. We refer more details of DPO derivation in Appendix A.2.
@@ 2.2 Motivation: Revisiting xPOs
Since DPO is proposed, there is huge interest in developing and improving DPO, leading to numerous xPOs. Different xPOs can be motivated by theoretical concerns like relaxing or extending preference distribution assumptions in IPO and MallowsPO, human aware loss function in KTO, or from practical aspects like reference model-free alternatives, like CPO, ORPO and SimPO. We provide an non-exhaustive list in Table 8 in Appendix B for the ease of revisiting and comparison.
@@
Despite their differing motivations, xPOs share a primary objective to optimize. We thus take the loss objectives as the first class citizen, and mathematically understand the parts that are commonly adopted or differ in xPOs. Before going into detailed categorization, we want to first argue that, in existing preference optimization literature, there is a lack of comprehensive studies on revisiting and examining DPO variants in their mathmatical objectives. As a consequence, some papers may have implicitly proposed some designs for improvement and even didn’t highlight it. As a motivating example, we revisit ORPO objective, which proposes to maximize an odd ratio difference (for an event A with probability $p$, the odds ratio is defined as $p/(1-p)$) between the winning and losing answers:
@@
$\mathcal{L}_{\mathrm{ORPO}}\left(\pi_{\theta}\right)=-\mathbb{E}[\underbrace{\log p_{\theta}(y_{w}|x)}_{\mathcal{L}_{\text{SFT}}\left(\pi_{\theta}\right)}+\underbrace{\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right)}_{\lambda\cdot\mathcal{L}_{\text{PO}}\left(\pi_{\theta}\right)}],\vspace{-7pt}$ (3)
@@
in which the expectation is for $\left(x,y_{w},y_{l}\right)\sim\mathcal{D}$, and $p_{\theta}(y|x)=\exp\left(\frac{1}{|y|}\log\pi_{\theta}(y|x)\right)$. Rewriting terms in $\mathcal{L}_{\text{PO}}\left(\pi_{\theta}\right)$, we could derive a upper bound as (see proof in Appendix D.1):
@@
$\mathcal{L}_{\mathrm{PO}}\left(\pi_{\theta}\right)\leq-\log\sigma(\frac{1}{1-p_{\theta}(y_{l}|x)}\underbrace{\left(\frac{1}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{1}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)\right)}_{\Delta_{\theta}}):=\bar{\mathcal{L}}_{\text{PO}}\left(\pi_{\theta}\right),\vspace{-5pt}$ (4)
if assuming $\Delta_{\theta}>0$ for all $x$. The upper bound ¯$\bar{\mathcal{L}}_{\text{PO}}$ is sharp, as ¯$\bar{\mathcal{L}}_{\text{PO}}-\mathcal{L}_{\text{PO}}=\mathcal{O}(\Delta_{\theta})^{2}$; thus minimizing ORPO loss could be understood as reference-model free DPO with length normalization (namely $1/|y_{w}|$ and $1/|y_{l}|$, see more detailed explanation in Section 3) and a contextual dependent $\beta(x)=1/(1-p_{\theta}(y_{l}|x))$. Length normalization is one of the key ideas adopted in SimPO:
$\mathcal{L}_{\mathrm{SimPO}}\left(\pi_{\theta};\gamma\right):=-\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right],$ (5)
@@
which is evident in SimPO objective, though proposed after ORPO. This connection is however unaware by the literature. This thus calls for a comprehensive analysis of the contributed elements in different xPOs so far, as many methods may overlap in contributed directions without awareness, and bringing this out right away could possibly prevent repetitive work or efforts in the future.
@@
Following similar analysis of different representative XPO methods for pairwise preferences, including DPO (Rafailov et al., 2023), IPO (Azar et al., 2024), CPO (Xu et al., 2024), GPO (Tang et al., 2024), RSO (Liu et al., 2023), ODPO (Amini et al., 2024), ORPO (Hong et al., 2024), MallowsPO (Chen et al., 2024a), SimPO (Meng et al., 2024), we come up with seven broad categories, which is able to explain most popular DPO variants in the literature, as in Table 1, This provides a straightforward illustration of the main ideas and connections of existing methods. The meanings and details of the categories are elaborated in Section 3.
@@
MallowsPO $\times$ logistic $\times$ SFT $\checkmark$ $\times$ $\times$
@@
Table 1: Mapping of xPOs with mathematically orthogonal components (see more details in Appendix C) and validation results of their effectiveness by the downstream task evaluations.
@@ 3.1 Component Descriptions
We first explain in detail about the components we categorized, after which we propose a generic framework, which we name as RainbowPO, to combine these components.
@@
The literature has noticed a verbosity issue of DPO aligned models, as the aligned model may generate answers significantly longer than both preferred and rejected answers (Park et al., 2024). This also could lead to an inflated win wate when evaluating model performance, as LLM-as-a-judge can be susceptible to length bias: Wang et al. (2023) has noticed that when evaluating 13B parameter models in head-to-head comparisons with the Davinci-003 model, win rates have a strong correlation (0.96) with the average number of unique tokens in the model’s response. To address this issue, one promising direction noticed in the literature is to incorporate explicit length penalties, like in R-DPO (Park et al., 2024) and SimPO (Meng et al., 2024):
@@
in which $r_{\theta}(x,y)=\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$ is the implicit reward model (Rafailov et al., 2023). From an optimization perspective, maximization with respect to $r^{\text{LN}}_{\theta}(x,y)$ is equivalent to $r^{\text{LR}}_{\theta}(x,y)$ with a specific $\alpha$ (might be prompt $x$ dependent). Directly applied to DPO, the resulting objective yields:
$\mathcal{L}_{\mathrm{LN-DPO}}\left(\pi_{\theta};\pi_{\text{ref}}\right):=-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}{\mathbb{E}}\log\sigma\left(\frac{\beta}{|y_{w}|}\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\text{ref}}\left(y_{w}\mid x\right)}-\frac{\beta}{|y_{l}|}\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\text{ref}}\left(y_{l}\mid x\right)}\right).$ (7)
Why length normalization could help prevent the verbosity issues can be explained through examining the gradient of the loss respectively: $\nabla_{\theta}\mathcal{L}_{\mathrm{LN-DPO}}\left(\pi_{\theta};\pi_{\text{ref}}\right)=$
$-\beta\mathbb{E}\left[\sigma\left(r^{\text{LN}}_{\theta}\left(x,y_{l}\right)-r^{\text{LN}}_{\theta}\left(x,y_{w}\right)\right)\left(\frac{1}{|y_{w}|}\nabla_{\theta}\log\pi_{\theta}\left(y_{w}\mid x\right)-\frac{1}{|y_{l}|}\nabla_{\theta}\log\pi_{\theta}\left(y_{l}\mid x\right)\right)\right],$ (8)
thus the gradient of length normalized DPO can be understood as taking a discount factor $\frac{1}{|y_{w}|}$ of the length for longer sequence. We also empirically justify the effectiveness of length normalization by comparing to the vanilla DPO trained models, and witness the consistent smaller average length, independent of the regularization constant $\beta$. See results of average length in Section 4.
@@
SLiC-HF (Zhao et al., 2023) and GPO (Tang et al., 2024) both realized that the DPO objective could be understood as taking $f$ (we refer this as link function) as $-\log\sigma(\cdot)$ in:
$\mathcal{L}_{\text{GPO}}=\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}{\mathbb{E}}\left[f\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right],$ (9)
thus unifying DPO, IPO, SLiC (without SFT loss) as instances of taking $f$ as logistic $-\log\sigma(\cdot)$, square $(x-1/2)^{2}$ or hinge function $\max(0,\delta-x)$ for some margin $\delta>0$ (see more details in Appendix C), separately. For identifying the best link function, we did an exclusive parameter search for DPO and IPO separately, however we found DPO, i.e. adopting $-\log\sigma(\cdot)$ as the link function, empirically performs better than IPO evaluated by AlpacaEval2, despite the weaker theoretical assumption of preferences in IPO. Thus we stick to the link function to be $-\log\sigma(\cdot)$ in this paper.
@@
which takes the losing prompt in a home advantage when $\gamma>0$. SimPO shows the effectiveness of this margin under the reference-free setup; however, when we adopt the margin for vanilla DPO (i.e. with the reference policy) with the optimal $\beta$, we do not witness an increase of the performance when adjusting the margin, either further adopting DPO with length normalization or not. In Figure 1(a), the performance steadily decreases when increasing the margin $\gamma$ in DPO, i.e. we optimize:
$\mathcal{L}_{\mathrm{DPO+}}\left(\pi_{\theta};\pi_{\mathrm{ref}},\gamma\right):=-\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\text{ref }}\left(y_{w}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\text{ref }}\left(y_{l}\mid x\right)}-\gamma\right)\right],$ (11)
This questions the true explanation about the effectiveness of the margin term in SimPO. We provide the answer as understanding margin as a reference policy right in the next paragraph.
@@
DPO takes the SFT policy as the reference policy motivated by the standard RLHF pipeline. However, recently proposed methods like CPO, ORPO, and SimPO (Xu et al., 2024; Hong et al., 2024; Meng et al., 2024) all suggested a reference-free objective could yield the same or even better performance. CPO and ORPO further utilized an extra SFT loss to force regularization, while for SimPO, such regularization is not enforced. Given our prior examination that home advantage can hardly improve over DPO, we argue that the margin term in SimPO loss Eq. 5 should be understood as a term for “reference policy” instead of the “home advantage”.
@@
Concretely, we could hypothesize that there exists a “good policy” $\pi_{\gamma}$ such that, for each prompt and preference pairs in the dataset, the normalized log likelihood ratio of preferred response to non-preferred response is a positive constant, which we denote as $\pi_{\gamma}$. We assume that $\pi_{\gamma}$’s normalized implicit reward model is perfect at in-distribution pairwise classification and yields $\frac{\pi_{\gamma}\left(y_{w}\mid x\right)^{1/|y_{w}|}}{\pi_{\gamma}\left(y_{l}\mid x\right)^{1/|y_{l}|}}=\exp(\gamma)$ for any $x$. If so, the loss of SimPO (defined as in Equation 5) could be rewritten as: